# The algebraic and geometric classification of

### v1

@@ 1.3.5. Compatible pre-Lie algebras defined on $\mathcal{C}_{05}^{\alpha\neq 0,\frac{1}{2},1}$
  • If $\alpha_{2}=0$, we obtain the parametric family:
    $\mathcal{C}_{25}^{\alpha,\beta,\gamma}:=\left\{\begin{tabular}[]{ll}$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\alpha e_{2},$\\ $e_{1}\ast e_{1}=\beta e_{1},$&$e_{1}\ast e_{2}=\gamma e_{2}.$\end{tabular}\right.$
@@
1.3.6. Compatible pre-Lie algebras defined on $\mathcal{C}_{05}^{\frac{1}{2}}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{05}^{\frac{1}{2}},\mathcal{C}_{05}^{\frac{1}{2}})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{\frac{1}{2}}$ are of the form:
@@ 1.3.6. Compatible pre-Lie algebras defined on $\mathcal{C}_{05}^{\frac{1}{2}}$
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\frac{1}{2}e_{2},$\\ $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2},$&$e_{1}\ast e_{2}=\alpha_{3}e_{2}.$\end{tabular}\right.$
@@
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\frac{1}{2}e_{2},$&\\ $e_{1}\ast e_{1}=2\alpha_{4}e_{1},$&$e_{1}\ast e_{2}=\alpha_{4}e_{2},$&$e_{2}\ast e_{2}=\alpha_{5}e_{1}.$\end{tabular}\right.$
@@
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\frac{1}{2}e_{2},$&&\\ $e_{1}\ast e_{1}=2\alpha_{6}e_{1},$&$e_{1}\ast e_{2}=\alpha_{7}e_{1}+\alpha_{6}e_{2},$&$e_{2}\ast e_{1}=2\alpha_{7}e_{1}$&$e_{2}\ast e_{2}=\alpha_{8}e_{1}+\alpha_{7}e_{2}.$\end{tabular}\right.$
We may assume $\alpha_{5}\alpha_{7}\neq 0$.
If the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{\frac{1}{2}}$ is of the first form, then we obtain the algebras $\mathcal{C}_{24}^{\frac{1}{2},\beta,\gamma}$ and $\mathcal{C}_{25}^{\frac{1}{2},\beta,\gamma}$.
Assume now that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{\frac{1}{2}}$ is of the second form. Then, choose $\xi={\alpha^{-\frac{1}{2}}_{5}}$ and obtain the parametric family:
$\mathcal{C}_{26}^{\alpha}:=\left\{\begin{tabular}[]{lll}$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\frac{1}{2}e_{2},$&\\ $e_{1}\ast e_{1}=2\alpha e_{1},$&$e_{1}\ast e_{2}=\alpha e_{2},$&$e_{2}\ast e_{2}=e_{1}.$\end{tabular}\right.$
@@
Finally, assume that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{\frac{1}{2}}$ is of the third form. Then, choose $\xi={\alpha^{-1}_{7}}$ and obtain the parametric family:
$\mathcal{C}_{27}^{\alpha,\beta}:=\left\{\begin{tabular}[]{llll}$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\frac{1}{2}e_{2},$&&\\ $e_{1}\ast e_{1}=2\alpha e_{1},$&$e_{1}\ast e_{2}=e_{1}+\alpha e_{2},$&$e_{2}\ast e_{1}=2e_{1}$&$e_{2}\ast e_{2}=\beta e_{1}+e_{2}.$\end{tabular}\right.$
@@
1.3.7. Compatible pre-Lie algebras defined on $\mathcal{C}_{05}^{1}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{05}^{1},\mathcal{C}_{05}^{1})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{1}$ are of the form:
@@ 1.3.7. Compatible pre-Lie algebras defined on $\mathcal{C}_{05}^{1}$
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$\\ $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2},$&$e_{1}\ast e_{2}=\alpha_{3}e_{2}.$\end{tabular}\right.$
@@
    {$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$&&\\ $e_{1}\ast e_{1}=\alpha_{4}e_{1},$&$e_{1}\ast e_{2}=\alpha_{4}e_{2},$&$e_{2}\ast e_{1}=\alpha_{5}e_{1},$&$e_{2}\ast e_{2}=\alpha_{5}e_{2}.$\end{tabular}\right.$
We may assume $\alpha_{5}\neq 0$.
First, suppose that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{1}$ is of the first form. Then we have the following cases:
@@
  • If $\alpha_{1}=\alpha_{3}$, then we will consider the following two cases.
@@
Assume now that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{1}$ is of the second form. Then, choose $\xi={\alpha^{-1}_{5}},\nu=-{\alpha_{4}}{\alpha^{-1}_{5}}$ and obtain the algebra:
$\mathcal{C}_{28}:=\left\{\begin{tabular}[]{ll}$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=e_{2},$\\ $e_{2}\ast e_{1}=e_{1},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
1.3.8. Compatible pre-Lie algebras defined on $\mathcal{C}_{05}^{0}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{05}^{0},\mathcal{C}_{05}^{0})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{0}$ are of the form:
@@ 1.3.8. Compatible pre-Lie algebras defined on $\mathcal{C}_{05}^{0}$
    {$e_{1}\cdot e_{1}=e_{1},$&\\ $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2},$&$e_{1}\ast e_{2}=\alpha_{3}e_{2}.$\end{tabular}\right.$
@@
    {$e_{1}\cdot e_{1}=e_{1},$&\\ $e_{1}\ast e_{1}=\alpha_{4}e_{1},$&$e_{2}\ast e_{2}=\alpha_{5}e_{2}.$\end{tabular}\right.$
@@
    {$e_{1}\cdot e_{1}=e_{1},$&&&\\ $e_{1}\ast e_{1}=\alpha_{6}e_{1}+\alpha_{7}e_{2},$&$e_{1}\ast e_{2}=\alpha_{8}e_{1},$&$e_{2}\ast e_{1}=\alpha_{8}e_{1},$&$e_{2}\ast e_{2}=\alpha_{8}e_{2}.$\end{tabular}\right.$
We may assume $\alpha_{5}\alpha_{8}\neq 0$.
If the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{0}$ is of the first form, then we obtain the algebras $\mathcal{C}_{24}^{0,\beta,\gamma}$ and $\mathcal{C}_{25}^{0,\beta,\gamma}$.
Assume now that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{0}$ are of the second form. Then, choose $\xi={\alpha^{-1}_{5}}$ and obtain the parametric family:
$\mathcal{C}_{29}^{\alpha}:=\left\{\begin{tabular}[]{ll}$e_{1}\cdot e_{1}=e_{1},$&\\ $e_{1}\ast e_{1}=\alpha e_{1},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
@@
Finally, assume that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{0}$ is of the third form. Then, choose $\xi={\alpha^{-1}_{8}}$ and obtain the parametric family:
$\mathcal{C}_{30}^{\alpha,\beta}:=\left\{\begin{tabular}[]{llll}$e_{1}\cdot e_{1}=e_{1},$&&&\\ $e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},$&$e_{1}\ast e_{2}=e_{1},$&$e_{2}\ast e_{1}=e_{1},$&$e_{2}\ast e_{2}=e_{2}.$\end{tabular}\right.$
@@
1.3.9. Compatible pre-Lie algebras defined on $\mathcal{C}_{06}^{\alpha\neq 0,1,2}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{06}^{\alpha},\mathcal{C}_{06}^{\alpha})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{\alpha}$ are of the form:
=e1e1e1, =e1e2αe2, =e2e1e2,
=e1e1+α1e1α2e2, =e1e2α3e2, =e2e1α1e2.
[…]
Then we have the following cases:
  • If $\alpha_{2}\neq 0$, then choose $\xi=\alpha_{2}$ and obtain the parametric family:
    $\mathcal{C}_{31}^{\alpha,\beta,\gamma}:=\left\{\begin{tabular}[]{lll}$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\alpha e_{2},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\beta e_{1}+e_{2},$&$e_{1}\ast e_{2}=\gamma e_{2},$&$e_{2}\ast e_{1}=\beta e_{2}.$\end{tabular}\right.$
@@ 1.3.9. Compatible pre-Lie algebras defined on $\mathcal{C}_{06}^{\alpha\neq 0,1,2}$
  • If $\alpha_{2}=0$, we obtain the parametric family:
    $\mathcal{C}_{32}^{\alpha,\beta,\gamma}:=\left\{\begin{tabular}[]{lll}$e_{1}\cdot e_{1}=e_{1},$&$e_{1}\cdot e_{2}=\alpha e_{2},$&$e_{2}\cdot e_{1}=e_{2},$\\ $e_{1}\ast e_{1}=\beta e_{1},$&$e_{1}\ast e_{2}=\gamma e_{2},$&$e_{2}\ast e_{1}=\beta e_{2}.$\end{tabular}\right.$
@@
1.3.10. Compatible pre-Lie algebras defined on $\mathcal{C}_{06}^{0}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{06}^{0},\mathcal{C}_{06}^{0})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{0}$ are of the forms:

### v2

@@ 2.3.5 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{05}^{\alpha\not=0,\frac{1}{2},1}}$
  • If $\alpha_{2}=0$, we obtain the parametric family
    $\mathcal{C}_{25}^{\alpha,\beta,\gamma}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=\alpha e_{2},\quad e_{1}\ast e_{1}=\beta e_{1},\quad e_{1}\ast e_{2}=\gamma e_{2}.$
@@
2.3.6 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{05}^{\frac{1}{2}}}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}\bigl{(}\mathcal{C}_{05}^{\frac{1}{2}},\mathcal{C}_{05}^{\frac{1}{2}}\bigr{)}$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{\frac{1}{2}}$ are of the following forms:
@@ 2.3.6 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{05}^{\frac{1}{2}}}$
    $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\cdot e_{2}=\frac{1}{2}e_{2}$, $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2}$, $e_{1}\ast e_{2}=\alpha_{3}e_{2}$,
@@
    $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\cdot e_{2}=\frac{1}{2}e_{2}$, $e_{1}\ast e_{1}=2\alpha_{4}e_{1}$, $e_{1}\ast e_{2}=\alpha_{4}e_{2}$, $e_{2}\ast e_{2}=\alpha_{5}e_{1}$,
@@
    $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\cdot e_{2}=\frac{1}{2}e_{2}$, $e_{1}\ast e_{1}=2\alpha_{6}e_{1}$, $e_{1}\ast e_{2}=\alpha_{7}e_{1}+\alpha_{6}e_{2}$, $e_{2}\ast e_{1}=2\alpha_{7}e_{1}$, $e_{2}\ast e_{2}=\alpha_{8}e_{1}+\alpha_{7}e_{2}$.
We may assume $\alpha_{5}\alpha_{7}\neq 0$. If the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{\frac{1}{2}}$ is of the first form, then we obtain the algebras $\mathcal{C}_{24}^{\frac{1}{2},\beta,\gamma}$ and $\mathcal{C}_{25}^{\frac{1}{2},\beta,\gamma}$.
Assume now that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{\frac{1}{2}}$ is of the second form. Then, choose $\xi={\alpha^{-\frac{1}{2}}_{5}}$ and obtain the parametric family
$\mathcal{C}_{26}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=\frac{1}{2}e_{2},\quad e_{1}\ast e_{1}=2\alpha e_{1},\quad e_{1}\ast e_{2}=\alpha e_{2},\quad e_{2}\ast e_{2}=e_{1}.$
@@
Finally, assume that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{\frac{1}{2}}$ is of the third form. Then, choose $\xi={\alpha^{-1}_{7}}$ and obtain the parametric family
$\mathcal{C}_{27}^{\alpha,\beta}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=\tfrac{1}{2}e_{2},\quad e_{1}\ast e_{1}=2\alpha e_{1},$
$\ e_{1}\ast e_{2}=e_{1}+\alpha e_{2},\quad e_{2}\ast e_{1}=2e_{1},\quad e_{2}\ast e_{2}=\beta e_{1}+e_{2}.$
@@
2.3.7 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{05}^{1}}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}\bigl{(}\mathcal{C}_{05}^{1},\mathcal{C}_{05}^{1}\bigr{)}$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{1}$ are of the following forms:
@@ 2.3.7 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{05}^{1}}$
    $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\cdot e_{2}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2}$, $e_{1}\ast e_{2}=\alpha_{3}e_{2}$,
@@
    $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\cdot e_{2}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{4}e_{1}$, $e_{1}\ast e_{2}=\alpha_{4}e_{2}$, $e_{2}\ast e_{1}=\alpha_{5}e_{1}$, $e_{2}\ast e_{2}=\alpha_{5}e_{2}$.
We may assume $\alpha_{5}\neq 0$. First, suppose that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{1}$ is of the first form. Then we have the following cases:
@@
  • If $\alpha_{1}=\alpha_{3}$, then we will consider the following two cases:
@@
Assume now that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{1}$ is of the second form. Then, choose $\xi={\alpha^{-1}_{5}}$, $\nu=-{\alpha_{4}}{\alpha^{-1}_{5}}$ and obtain the algebra
$\mathcal{C}_{28}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=e_{2},\quad e_{2}\ast e_{1}=e_{1},\quad e_{2}\ast e_{2}=e_{2}.$
2.3.8 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{05}^{0}}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}\bigl{(}\mathcal{C}_{05}^{0},\mathcal{C}_{05}^{0}\bigr{)}$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{0}$ are of the following forms:
@@ 2.3.8 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{05}^{0}}$
    $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2}$, $e_{1}\ast e_{2}=\alpha_{3}e_{2}$,
@@
    $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\ast e_{1}=\alpha_{4}e_{1}$, $e_{2}\ast e_{2}=\alpha_{5}e_{2}$,
@@
    $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\ast e_{1}=\alpha_{6}e_{1}+\alpha_{7}e_{2}$, $e_{1}\ast e_{2}=\alpha_{8}e_{1}$, $e_{2}\ast e_{1}=\alpha_{8}e_{1}$, $e_{2}\ast e_{2}=\alpha_{8}e_{2}$.
We may assume $\alpha_{5}\alpha_{8}\neq 0$. If the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{0}$ is of the first form, then we obtain the algebras $\mathcal{C}_{24}^{0,\beta,\gamma}$ and $\mathcal{C}_{25}^{0,\beta,\gamma}$.
Assume now that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{0}$ are of the second form. Then, choose $\xi={\alpha^{-1}_{5}}$ and obtain the parametric family
$\mathcal{C}_{29}^{\alpha}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\ast e_{1}=\alpha e_{1},\quad e_{2}\ast e_{2}=e_{2}.$
@@
Finally, assume that the compatible pre-Lie algebra structures defined on $\mathcal{C}_{05}^{0}$ is of the third form. Then, choose $\xi={\alpha^{-1}_{8}}$ and obtain the parametric family
$\mathcal{C}_{30}^{\alpha,\beta}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\ast e_{1}=\alpha e_{1}+\beta e_{2},\quad e_{1}\ast e_{2}=e_{1},\quad e_{2}\ast e_{1}=e_{1},\quad e_{2}\ast e_{2}=e_{2}.$
@@
2.3.9 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{06}^{\alpha\neq 0,1,2}}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}(\mathcal{C}_{06}^{\alpha},\mathcal{C}_{06}^{\alpha})$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{\alpha}$ are of the form $e_{1}\cdot e_{1}=e_{1}$, $e_{1}\cdot e_{2}=\alpha e_{2}$, $e_{2}\cdot e_{1}=e_{2}$, $e_{1}\ast e_{1}=\alpha_{1}e_{1}+\alpha_{2}e_{2}$, $e_{1}\ast e_{2}=\alpha_{3}e_{2}$, $e_{2}\ast e_{1}=\alpha_{1}e_{2}$. Then we have the following cases:
  • If $\alpha_{2}\neq 0$, then choose $\xi=\alpha_{2}$ and obtain the parametric family
    $\mathcal{C}_{31}^{\alpha,\beta,\gamma}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=\alpha e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
    $\ e_{1}\ast e_{1}=\beta e_{1}+e_{2},\quad e_{1}\ast e_{2}=\gamma e_{2},\quad e_{2}\ast e_{1}=\beta e_{2}.$
@@ 2.3.9 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{06}^{\alpha\neq 0,1,2}}$
  • If $\alpha_{2}=0$, we obtain the parametric family
    $\mathcal{C}_{32}^{\alpha,\beta,\gamma}\colon\ e_{1}\cdot e_{1}=e_{1},\quad e_{1}\cdot e_{2}=\alpha e_{2},\quad e_{2}\cdot e_{1}=e_{2},$
    $\ e_{1}\ast e_{1}=\beta e_{1},\quad e_{1}\ast e_{2}=\gamma e_{2},\quad e_{2}\ast e_{1}=\beta e_{2}.$
@@
2.3.10 Compatible pre-Lie algebras defined on $\boldsymbol{\mathcal{C}_{06}^{0}}$
From the computation of $\mathrm{Z}_{\mathrm{CPL}}^{2}\bigl{(}\mathcal{C}_{06}^{0},\mathcal{C}_{06}^{0}\bigr{)}$, the compatible pre-Lie algebra structures defined on $\mathcal{C}_{06}^{0}$ are of the following forms: